7.2 × 10-3 kg⋅m/s.
1
Expert's answer
2019-09-12T10:36:43-0400

This is a partiicular value of momentum of some object:


"p=mv=7.2\cdot 10^{-3}\frac{kg\cdot m}{s}"

Where m is its mass and v is its speed.